OpenStudy (anonymous):

Riemann Sum: midpoint -9x^2-5y m=n=2 [0,2]x[0,4]

OpenStudy (anonymous):

so the points would be (1/2, 1) (3/2, 1) (1/2, 3) (3/2, 3)

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

area = 2

OpenStudy (anonymous):

2 * ( -(29/4) -(101/4) -(69/4) -(141/4) )

OpenStudy (anonymous):

= -170 but it's area so it would be 170

OpenStudy (anonymous):

just want to make sure that's right cuz i did the lower left hand and got 76 but it was wrong

OpenStudy (anonymous):

lower left would be (0,0) (0,2) (1,0) (1,2) -10 * 2 -10 * 2 -19 * 2 answer 76?

OpenStudy (anonymous):

I don't follow your question whatsoever. The equation is \(-9x^2-5y=0\)? What are \(m,n\)? What's \(\left[0,2\right]\times\left[0,4\right]\)?

OpenStudy (anonymous):

it's a reimann sum for double integral (-9 x^2 - 5y) m the how many squares for the x axis n is how many squares to segment the y axis

OpenStudy (anonymous):

[0,2] = \[0 \le x \le2\]

OpenStudy (anonymous):

[0,4] = \[0 \le y \le 4\]

OpenStudy (anonymous):

cross product giving you the height

OpenStudy (anonymous):

Gotcha.

OpenStudy (anonymous):

here's a pic of the original problem

OpenStudy (anonymous):

hrm

OpenStudy (dumbcow):

looks right...i get an exact answer of -176 so for midpoint you are on right track

OpenStudy (anonymous):

ah k hopefully it takes! :0

OpenStudy (anonymous):

here goes

OpenStudy (anonymous):

it would be positive right?

OpenStudy (dumbcow):

oh yeah i guess if we are assuming positive volume

OpenStudy (anonymous):

er 176 didn't work

OpenStudy (anonymous):

and 76 didn't work for left hand

OpenStudy (dumbcow):

oh sorry thats the exact answer, you were corroct with 170 as midpoint estimate

OpenStudy (dumbcow):

correct*

OpenStudy (anonymous):

whew had 3 chances, last chance worked -170 was the answer

OpenStudy (anonymous):

do you mind double checking my lower left hand answer? :)

OpenStudy (anonymous):

pretty sure it is -76 done it about 4 times

OpenStudy (dumbcow):

yes i get -76 as well :)